OpenStudy (anonymous):

solve by substitution. x/3 + y/2 = 4 x/2 - y = -1

OpenStudy (anonymous):

please help guys ! =(

OpenStudy (dumbcow):

First get y by itself in bottom equation -> y = 1+x/2 Then replace this for y in top equation ->x/3 + (1+x/2)/2 = 4 solve for x by combining fractions using common denominators...

OpenStudy (anonymous):

and can you help me with the combining fractions part?

OpenStudy (dumbcow):

1+x/2 needs denominator of 2 -> 2/2 + x/2 = (x+2)/2

OpenStudy (dumbcow):

(x+2)/2 * 1/2 = (x+2)/4 with me?

OpenStudy (anonymous):

yes, now what?

OpenStudy (dumbcow):

Then x/3 + (x+2)/4 = 4 to add these we need denominator of 12 ->(x/3)*(4/4) + (x+2)/4*(3/3) = 4x/12 + 3(x+2)/12 = (4x + 3x + 6)/12

OpenStudy (dumbcow):

(7x+6)/12 = 4 multiply by 12 on both sides ->7x + 6 = 48 x = 6

OpenStudy (dumbcow):

did you follow that?

OpenStudy (anonymous):

yes thank you so much. now how do i figure out y?

OpenStudy (dumbcow):

plug in 6 for x into that equation above remember whn we solved for y initially ->y = 1+x/2 replace x with 6 ->y = 1+(6/2) = 1+3=4